Calculate the minimum speed would blood need to travel

through a small blood vessel with a radius of 1 mm before the flow

was turbulent?

Where the viscosity of the blood is ηblood = 2.7× 10−3 Pa s and its density is ρblood = 1050 kg m−3



1
Expert's answer
2021-03-21T11:25:34-0400

For flow in a pipe of diameter "D = 2r = 2mm = 2\times 10^{-3}m" laminar flow occurs when "Re_D= 2300" (upper boundary), where "Re_D" is the Reynolds number:


"Re_D = \dfrac{\rho u_{max} D}{\mu} = 2300"

where "\rho = 1050kg/m^3" is the density of the blood, and "\mu = 2.7\times 10^{-3}Pa\cdot s" is the dynamic viscosity of the blood, and "u_{max}" is the maximum speed. Expressing "u_{max}", obtain:


"u_{max} = \dfrac{2300\mu}{\rho D}\\
u_{max} = \dfrac{2300\cdot 2.7\times 10^{-3}}{1050 \cdot 2\times 10^{-3}} \approx 4.1m/s"

Answer. 4.1 m/s.